Find a vector function r(t) = <x(t),y(t),z(t)>

where r(t) is continuous everywhere except t=2


lim(t=2)r(t)= <1,0,0>



1
Expert's answer
2021-03-01T07:15:50-0500

"y(t)=z(t)=t-2"

"x(t)=1+e^{-\frac{1}{|t-2|}}"


"r(t)=<1+e^{-\frac{1}{|t-2|}},t-2,t-2>"